# Inertial Magnetic SLAM Systems Using Low-Cost Sensors

### v1

@@ Abstract
Spatially inhomogeneous magnetic fields offer a valuable, non-visual information source for positioning. Among systems leveraging this, magnetic field-based simultaneous localization and mapping (SLAM)systems are particularly attractive because they can provide positioning information and build a magnetic field map on the fly. Moreover, they have bounded error within mapped regions. However, state-of-the-art methods typically require low-drift odometry data provided by visual odometry or a wheel encoder, etc. This is because these systems need to minimize/reduce positioning errors while exploring, which happens when they are in unmapped regions. To address these limitations, this work proposes a loosely coupled and a tightly coupled inertial magnetic SLAM (IM-SLAM)system. The proposed systems use commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer. The use of non-visual data provides a significant advantage over visual-based systems, making it robust to low-visibility conditions. Both systems employ state-space representations, and magnetic field models on different scales. The difference lies in how they use a local and global magnetic field model. The loosely coupled system uses these models separately in two state-space models, while the tightly coupled system integrates them into one state-space model. Experiment results show that the tightly coupled IM-SLAMsystem achieves lower positioning errors than the loosely coupled system in most scenarios, with typical errors on the order of meters per 100 meters traveled. These results demonstrate the feasiblity of developing a full 3D IM-SLAMsystems using low-cost sensors and the potential of applying these systems in emergency response scenarios such as mine/fire rescue.
{IEEEkeywords}
@@ 1 Introduction
Positioning and navigation services are essential for various applications, including scientific research, emergency response, and military operations. These services have historically relied on the global navigation satellite system (GNSS). However, GNSS-denied environments, such as deep sea, indoor environments, and battlefields, where GNSS signals are unavailable or unreliable, pose significant challenges for accurate and reliable positioning. Therefore, developing robust and precise positioning systems that can operate effectively in GNSS-denied environments is of paramount importance.
@@
A common approach to address the challenges in positioning in GNSS-denied environments is to use additional sensors, such as cameras, Light Detection and Ranging (LiDAR)sensors, and inertial measurement units (IMUs), to complement the GNSS. For example, visual and LiDARSimultaneous localization and mapping (SLAM) systems [1, 2] have been widely used for indoor positioning. However, these systems have their own limitations. Visual SLAM systems are sensitive to lighting conditions and may fail in low-light or featureless environments [3], and LiDAR SLAM systems can be expensive and may function poorly in environments with limited geometric features [4]. Thus, it is necessary to exploit other types of environmental features to achieve robust indoor SLAM.
@@
Indoor magnetic fields have been recognized as a promising source of environmental features for SLAM [5]. Because of the ubiquitous presence of ferromagnetic materials in buildings, the indoor magnetic field exhibits spatial variations, which can be used as location fingerprints. As is shown in Fig. 1, the magnetic field strength inside a building can vary significantly over short distances, making it a useful feature for SLAMapplications. Moreover, magnetic field measurements can be obtained using low-cost magnetometers, which are commonly found in smartphones and other portable devices. This makes magnetic field SLAMsystems cost-effective and accessible.
@@
Previous works have shown that magnetic field SLAMsystems can achieve meter-level accuracy in indoor environments [6, 7, 8, 9, 10]. However, most of these systems rely on low-drift odometry data, such as that from wheel odometry or visual odometry, to provide motion constraints, or they assume planar 2D motion, which limits their availability and applicability in real-world scenarios. Furthermore, these systems use a single magnetometer, which makes SLAMmore challenging because a measurement of the magnetic field at a single point provides limited information about the spatial variations in the magnetic field. To address these issues, in a recent work [11], a magnetic field-aided inertial navigation system was developed. This system uses a low-cost IMUand a magnetometer array consisting of 30 magnetometers, which allows snapshots of magnetic field over a planar sensor board to be taken. This system can provide low-drift inertial navigation that has the potential to be used, instead of wheel odometry or visual odometry, in a magnetic-field SLAM system. Inspired by this work, we seek to answer the following research questions:
Figure 1: Illustration of the magnetic-field magnitude variations, represented by different colors, near the floor in a room of $12\times 12$ square meters.
@@
  • Can a 3D inertial magnetic SLAM (IM-SLAM) system be built and provide meter-level positioning using only a low-cost IMU, barometer, and magnetometer array?
@@
  • What is a suitable sensor fusion strategy? How does the performance of a tightly-coupled IM-SLAMsystem compare to a loosely-coupled IM-SLAMsystems?
@@
  • How does the barometer and the quality of the inertial sensor measurements affect the overall performance of the SLAM system?
@@
To that end, we propose two IM-SLAMsystems: a loosely coupled and a tightly coupled system. Different from previous magnetic field SLAMsystems, the proposed systems do not require odometric information from visual or wheel-encoder sensors. This makes the systems much more adaptable to various environments where traditional odometry might fail.
@@ 1.1 Related Work
Many methods for magnetic field-based indoor positioning and navigation have been proposed. They are generally classified into two categories: infrastructure-based methods [12] and infrastructure-free methods [13, 14]. The former generally requires some magnetic source to be installed in the environment, such as magnetic coils or magnets. The latter, on the other hand, exploits the ambient magnetic field as a location fingerprint. The infrastructure-free methods are more appealing because they are easier to deploy and maintain. Therefore, our focus is on the infrastructure-free methods.
The infrastructure-free methods can be further classified into two categories: fingerprinting-based methods[15, 16, 17, 18, 14, 19] and model-based methods [20, 21, 11, 13, 7, 22, 10, 6]. The fingerprinting-based methods are based upon the assumption that the magnetic fields at different locations are distinct. These methods generally rely on matching real-time magnetic field measurements to a pre-established magnetic field map to estimate location, or to a stored measurement history to recognize previously visited locations for loop closure. The model-based methods, on the other hand, seek to use a mathematical model to describe magnetic fields measurements. These methods can use the magnetic measurements to update their magnetic field model (build a magnetic field map) and estimate pose change as the device moves through the environment. There are pros and cons for both methods. The fingerprinting-based methods can achieve high accuracy if magnetic field matching is successful, but they require carefully designed feature extraction and matching algorithms. Furthermore, erroneous matching can lead to significant localization errors. In contrast, the model-based methods are easier to analyze from a theoretical perspective, but they can be sensitive to the choice of the magnetic field model and the tuning of the model parameters.
The proposed IM-SLAMsystems are model-based. They were made possible by the development in magnetic field odometry and previous works on magnetic field modeling. In [20], a polynomial model was proposed to locally approximate the magnetic field, which was later extended in [21] to estimate pose changes of a magnetometer array between two time instants. The very same polynomial model was also used in [11] to build a magnetic field aided inertial navigation system (MAINS), which effectively reduced the positioning drift of the inertial navigation system by 2 orders of magnitude. This work constitutes one of the building blocks of the proposed SLAM systems. The other building block comes from the work in global magnetic field modeling. Originally proposed in [23], the reduced rank Gaussian process (GP)model was shown to be effective and computationally efficient in modeling the indoor magnetic field on a global scale. The GP model was later used in [24, 6, 7] to build magnetic field SLAM systems.
@@ 1.2 Contributions
  • A loosely-coupled and tightly-coupled IM-SLAMsystems using low-cost sensors are proposed.
  • An extensive evaluation is conducted on real-world datasets, encompassing challenging indoor environments with multi-floor structures.
  • The implementation of the proposed systems and the datasets used for evaluation are made publicly available to provide a reproducible benchmark.
@@ 2 Preliminaries
As magnetic field modeling is essential for magnetic field SLAM, we briefly review two commonly used magnetic field models in this section; the first model is for modeling the magnetic field at local small scales, and the second is at global large scales.
Let us start with magnetostatics. Assuming that there is no time-varying electric field and current density in the region of interest, the magnetic field fulfills the curl-free condition [25, p. 180]
@@
Both of the two magnetic field models presented next are based on (2).
@@ 2.1 Local Magnetic Field Model
Here $c_{i,j,k}\in\mathbb{R}$ denotes the coefficient and $r\triangleq[r_{x}\;r_{y}\;r_{z}]^{\top}$. Furthermore, the coefficients of the resulting magnetic field model $M(r;c)$ from applying (2) are further constrained by the divergence-free condition [25, p. 180]
@@ 2.2 Global Magnetic Field Model
A global magnetic field model refers to a model that describes the magnetic field across the entire environment, such as a building or a large lab. The approximated GPmodel proposed in [23] allows for efficient regression with much less computational complexity than the full GPmodel. The development of the magnetic field model starts with modeling the scalar magnetic potential function $\phi(r)$ as a GP
@@
and
$\kappa_{\text{SE}}(r,r^{\prime})=\sigma_{\text{SE}}^{2}\text{exp}\left(-\frac{\lVert r-r^{\prime}\rVert^{2}}{2l^{2}_{\text{SE}}}\right).$ (7c)
Here $\sigma_{\text{lin.}}$ denotes the magnitude scale of the linear kernel $\kappa_{\text{lin.}}$, $\sigma_{\text{SE}}$ and $l_{\text{SE}}$ denote the magnitude scale and the characteristic length scale of the square exponential kernel $\kappa_{\text{SE}}$, respectively. To reduce the computation complexity, the square exponential kernel is approximated by
@@
where $m$ is the number of basic functions and $S_{\text{SE}}(\cdot)$ denotes the spectral density of $\kappa_{\text{SE}}$. Furthermore, $\psi_{j}(\cdot)$ and $\lambda_{j}$ denote the $j$-th eigenfunction and eigenvalue of the Laplace operator in the domain of a cuboid ($L_{x},L_{x}]\times[-L_{y},L_{y}]\times[-L_{z},L_{z}]$) with Dirichlet boundary conditions, respectively. They are given by
@@
respectively, where $n_{j,d}\in\mathbb{Z}^{+}$. Combining (7a) and (8a) and relating to the weight space view [26, Ch. 2], the scalar potential function is then approximated by
@@
$\Psi(r)=[r^{\top}\;\psi_{1}(r)\;\psi_{2}(r)\;\cdots\;\psi_{m}(r)]$ (9b)
$\eta\sim\mathcal{N}(0,\Lambda)$ (9c)
@@
$\Lambda=\text{diag}\left(\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},S_{\text{SE}}(\lambda_{1}),S_{\text{SE}}(\lambda_{2}),\cdots,S_{\text{SE}}(\lambda_{m})\right).$ (9d)
@@
2.3 Discusssion
The polynomial model presented in Section 2.1 is well-suited for local modeling, due to its simplicity and efficiency. The first-degree polynomial model was used in [11] to model the magnetic field within the small area covered by the magnetometer array at each time step (see Fig. 2). In this model, both the spatial coordinates and the magnetic field are expressed in the body frame of the magnetometer array.
On the other hand, the approximated GPmodel presented in Section 2.2 is more appropriate for global modeling, as the sinusoidal basis functions can capture the magnetic field variations across a larger area. It is more computationally demanding than the first-degree polynomial model. Typically, several hundreds to thousands of basis functions are used to represent the magnetic field across an area of several hundred square meters [7]. Since this model aims to describe the field over a global region, both the spatial coordinates and the magnetic field are expressed in the navigation frame.
Figure 2: Coordinate frames used by the local and global magnetic field models. The local magnetic field model at each time step aligns with the body frame of the magnetometer array, e.g., $\mathcal{B}_{1},\mathcal{B}_{2}$ and $\mathcal{B}_{3}$, while the global magnetic field model uses the navigation frame denoted by $O$ in the middle of the plot.
3 Magnetic Field-Based Positioning Systems
In this section, we briefly review the system models in two magnetic field-based positioning systems: the first is the magnetic field SLAM system proposed in [7], and the second is the MAINSproposed in [11]. These systems, illustrated in Fig. 3a and Fig. 3b, are the foundations of the proposed IM-SLAMsystems.
Figure 3: (a) An overview of the magnetic field SLAM system [7]. The system performs dead reckoning using visual odometry data and uses a global magnetic field model to register a single magnetometer’s measurements, which are used to correct position drift. (b) An overview of the MAINS. The system performs inertial navigation using IMUmeasurements and uses a local magnetic field model to register a magnetometer array’s measurements, which are used to correct velocity estimates.
@@ 3.1 The Magnetic Field SLAM System
The state-of-the-art magnetic field SLAM system [7] uses low-drift visual odometry and magnetic field measurements from a single magnetometer. The basic idea is to perform dead-reckoning using the visual odometry data and use magnetic field measurements to correct the accumulated position drift.
The state vector of the magnetic field SLAM at time step $k$ is defined as
$x_{k}^{\text{\tiny{SLAM}}}\triangleq[p_{k}^{\top}\;q_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (11)
where $p_{k}\in\mathbb{R}^{3}$ denotes the position, $q_{k}\in S^{3}$ denotes the orientation in quaternion representation, and $\eta_{k}\in\mathbb{R}^{3+m}$ denotes the coefficients of the approximated GPmagnetic field model.
@@
3.2 The Magnetic Field-Aided Inertial Navigation System
The MAINSproposed in [11] uses a low-cost IMUand magnetometer array (see Fig. 4) to provide low-drift inertial navigation. The basic idea is to use inertial navigation system (INS)to estimate poses, and magnetic field measurements to correct velocity estimates.
The state vector of MAINS at time step $k$ is defined as

### v2

@@ Abstract
Spatially inhomogeneous magnetic fields offer a valuable, non-visual information source for positioning. Among systems leveraging this, magnetic field-based simultaneous localization and mapping (SLAM) systems are particularly attractive. These systems execute positioning and magnetic field mapping tasks simultaneously, and they have bounded positioning error within previously visited regions. However, state-of-the-art magnetic-field SLAM methods typically require low-drift odometry data provided by visual odometry, a wheel encoder, or pedestrian dead-reckoning technology. To address this limitation, this work proposes loosely coupled and tightly coupled inertial magnetic SLAM (IM-SLAM) systems, which use only low-cost sensors: an inertial measurement unit (IMU), 30 magnetometers, and a barometer. Both systems are based on a magnetic-field-aided inertial navigation system (INS) and use error-state Kalman filters for state estimation. The key difference between the two systems is whether the navigation state estimation is done in one or two steps. These systems are evaluated in real-world indoor environments with multi-floor structures. The results of the experiment show that the tightly coupled inertial magnetic SLAM (IM-SLAM) system achieves lower positioning errors than the loosely coupled system in most scenarios, with typical errors on the order of meters per 100 meters traveled. These results demonstrate the feasibility of developing a full 3D IM-SLAM system using low-cost sensors. A potential application of the proposed systems is for the positioning of first responders.
keywords:
@@ 1 Introduction
Positioning and navigation services are essential for various applications, including scientific research, emergency response, and military operations. These services have historically relied on positioning information from global navigation satellite systems (GNSS). However, GNSS-denied environments, such as indoors or subsurface, where GNSS signals are unavailable or unreliable, pose significant challenges for accurate and reliable positioning. Therefore, developing robust and accurate positioning systems that can operate effectively in GNSS-denied environments is of paramount importance.
@@
A common approach to address the challenges of positioning in GNSS-denied environments is to use additional sensors, such as cameras, Light Detection and Ranging (LiDAR), and inertial measurement units (IMU s), to complement the GNSS. For example, visual and LiDAR SLAM systems [1, 2] have been widely used for indoor positioning. However, these systems have their own limitations. Visual SLAM systems are sensitive to lighting conditions and may fail in low-light or featureless environments [3], and LiDAR SLAM systems can be expensive and may function poorly in environments with limited geometric features [4]. Thus, it is necessary to exploit other types of environmental features to achieve robust indoor SLAM.
@@
Spatial variations of indoor magnetic fields have been recognized as a promising source of environmental features for SLAM [5]. Because of the ubiquitous presence of ferromagnetic materials in buildings, the indoor magnetic field exhibits spatial variations over short distances. Moreover, indoor magnetic fields can be measured using low-cost magnetometers, which are commonly found in smartphones and other portable devices. This makes magnetic field SLAM systems cost-effective and accessible.
@@
Previous works have shown that magnetic field SLAM systems can achieve meter-level accuracy in indoor environments [6, 7, 8, 10, 9]. However, most of these systems rely on low-drift odometry data, such as that from wheel odometry or visual odometry, to provide motion constraints, or they assume planar 2D motion, which limits their use in any type of practically useful positioning system. Furthermore, these systems use a single magnetometer, which makes SLAM more challenging because a measurement of the magnetic field at a single point provides limited information about the spatial variations in the magnetic field. To address these issues, in a recent work [11], a magnetic-field-aided inertial navigation system was developed. This system uses a low-cost IMU and an array of 30 magnetometers, which allows snapshots of the magnetic field to be taken. This system can provide low-drift inertial navigation that has the potential to be used, instead of wheel odometry or visual odometry, in a magnetic-field SLAM system. Inspired by this work, we seek to answer the following research questions:
@@
  • Can a 3D inertial magnetic SLAM (IM-SLAM) system be built using only a low-cost IMU, barometer, and magnetometer array?
@@
  • How does the performance of a tightly coupled IM-SLAM system compare to a loosely coupled IM-SLAM system?
@@
  • How do the barometer and the quality of the inertial sensor measurements affect the overall performance of the SLAM system?
@@
To that end, we propose two IM-SLAM systems: a loosely coupled and a tightly coupled system. Different from previous magnetic field SLAM systems, the proposed systems do not require odometric information from visual or wheel-encoder sensors. We evaluate the proposed systems on real-world datasets to answer the research questions.
@@ 1.1 Related Work
Many methods for magnetic-field-based indoor positioning have been proposed. They are generally classified into two categories: infrastructure-based methods [12] and infrastructure-free methods [13, 14]. The former generally requires some magnetic source to be installed in the environment, such as magnetic coils or magnets. The latter, on the other hand, exploits the ambient magnetic field as a location fingerprint. The infrastructure-free methods are more appealing because they are easier to deploy and maintain.
The infrastructure-free methods can be further classified into two categories: fingerprint-based methods[15, 16, 17, 18, 14, 19] and model-based methods [20, 21, 11, 13, 7, 22, 9, 6]. The fingerprint-based methods are based upon the assumption that the magnetic fields at different locations are distinct. These methods generally rely on matching magnetic field measurements to a pre-established magnetic field map to estimate position, or to a stored measurement history to recognize previously visited locations for loop closure. The model-based methods, on the other hand, seek to use a mathematical model to describe magnetic fields measurements. These methods can use the magnetic measurements to update their magnetic field model (build a magnetic field map) and estimate pose change as the device moves through the environment. There are pros and cons for both methods. The fingerprint-based methods can achieve high accuracy if magnetic field matching is successful, but they require carefully designed feature extraction and matching algorithms. Furthermore, erroneous matching can lead to significant positioning errors. In contrast, the model-based methods are easier to establish a theoretically sound framework for inference process, but they can be sensitive to the choice of the magnetic field model and the tuning of the model parameters.
The proposed IM-SLAM systems are model-based. They were made possible by developments in magnetic field odometry and previous work on magnetic field modeling. In [20], a polynomial model was proposed to locally approximate the magnetic field, which was later extended in [21] to estimate pose changes of a magnetometer array. This polynomial model was also used in [11] to build a magnetic field aided inertial navigation system (MAINS), which effectively reduced the positioning drift of the inertial navigation system by 2 orders of magnitude. This work [11] constitutes one of the building blocks of the proposed SLAM systems. The other building block comes from the work in global magnetic field modeling. Originally proposed in [23], the reduced rank Gaussian process (GP) model was shown to be effective and computationally efficient in modeling the indoor magnetic field on a global scale. The GP model was later used in [24, 6, 7] to build magnetic field SLAM systems.
@@ 1.2 Contributions
  • We propose a loosely coupled and tightly coupled 3D IM-SLAM system, which use only a low-cost IMU, a magnetometer array, and a barometer. These two system architectures serve as practical reference designs for practitioners.
  • We conduct an evaluation of the proposed systems in terms of their positioning performance on real-world datasets, including an indoor environment with multi-floor structures. The tightly coupled IM-SLAM system achieves a final positioning error of less than 5 meters over an approximately 200-meter trajectory.
  • We conduct ablation studies evaluating the impact of pose aiding for system initialization, barometric aiding for height stabilization, and inertial sensor quality. These analyses serve to clearly illustrate the overall performance bounds of the proposed systems.
@@ 2 Preliminaries
As magnetic field modeling is essential for magnetic field SLAM, we briefly review two commonly used magnetic field models: the first captures local, small-scale variations, while the second describes global, large-scale fields.
In the absence of time-varying electric fields and current density in the region of interest, the magnetic field fulfills the curl-free condition [25, p. 180]
@@
Both magnetic field models presented next are based on (2).
@@ 2.1 Local Magnetic Field Model
Here $c_{i,j,k}\in\mathbb{R}$ denotes the coefficient and $r\triangleq[r_{x}\;r_{y}\;r_{z}]^{\top}$ denotes the position. Furthermore, the coefficients of the resulting magnetic field model $M(r;c)$ from applying (2) are further constrained by the divergence-free condition [25, p. 180]
@@ 2.2 Global Magnetic Field Model
A global magnetic field model refers to a model that describes the magnetic field across the entire environment, such as a building or a large lab. The approximated GP model proposed in [23] allows for efficient regression with much less computational complexity than a full GP model. The development of the magnetic field model starts with modeling the scalar magnetic potential function $\phi(r)$ as a GP
@@
$\kappa_{\text{SE}}(r,r^{\prime})$ $=\sigma_{\text{SE}}^{2}\exp\left(-\frac{\lVert r-r^{\prime}\rVert^{2}}{2l^{2}_{\text{SE}}}\right).$ (7c)
Here $\mathcal{GP}\left(\mu,\kappa(x,x^{\prime})\right)$ denotes a collection of Gaussian random variables with mean $\mu$ and their covariance specified by $\kappa(x,x^{\prime})$. Furthermore, $\sigma_{\text{lin.}}$ denotes the magnitude scale of the linear kernel $\kappa_{\text{lin.}}$, $\sigma_{\text{SE}}$ and $l_{\text{SE}}$ denote the magnitude scale and the characteristic length scale of the square exponential kernel $\kappa_{\text{SE}}$, respectively. To reduce the computation complexity, the square exponential kernel is approximated by
@@
where $m$ is the number of basic functions and $S_{\text{SE}}(\cdot)$ denotes the spectral density of $\kappa_{\text{SE}}$. Furthermore, $\psi_{j}(\cdot)$ and $\lambda_{j}$ denote the $j$-th eigenfunction and eigenvalue of the Laplace operator in the domain of a cuboid ($L_{x},L_{x}]\times[-L_{y},L_{y}]\times[-L_{z},L_{z}]$) with Dirichlet boundary conditions. They are given by
@@
where $n_{j,d}\in\mathbb{Z}^{+}$. Combining (7a) and (8a) and relating to the weight space view [26, Ch. 2], the scalar potential function is then approximated by
@@
$\Psi(r)=[r^{\top}\;\psi_{1}(r)\;\psi_{2}(r)\;\cdots\;\psi_{m}(r)],\;\eta\sim\mathcal{N}(0,\Lambda)$ (9b)
@@
$\Lambda=\text{diag}\left(\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},S_{\text{SE}}(\lambda_{1}),S_{\text{SE}}(\lambda_{2}),\cdots,S_{\text{SE}}(\lambda_{m})\right).$ (9c)
@@
2.3 Discussion
The polynomial model presented in Section 2.1 is well-suited for local modeling, due to its simplicity and computational efficiency. The first-degree polynomial model was used in [11] to model the magnetic field within the small area covered by the magnetometer array at each time step. In this model, both the spatial coordinates and the magnetic field are typically expressed in the body frame of the magnetometer array.
On the other hand, the approximated GP model presented in Section 2.2 is appropriate for global modeling, as the sinusoidal basis functions can capture the magnetic field variations across a larger area. Typically, several hundreds to thousands of basis functions are used to represent the magnetic field across an area of several hundred square meters [7]. Therefore, this model is computationally costly. Since this model aims to describe the field over a global region, both the spatial coordinates and the magnetic field are typically expressed in the navigation frame.
3 Magnetic-Field-Based Positioning Systems
In this section, we briefly review the system models used in two magnetic field-based positioning systems. The first model is used by the magnetic field SLAM system in [7], and the second by MAINS in [11].
@@ 3.1 The Magnetic Field SLAM System
The state-of-the-art magnetic field SLAM system [7] uses low-drift odometry and magnetic field measurements from a single magnetometer. The state vector of the magnetic field SLAM at time step $k$ is defined as $x_{k}^{\text{\tiny{SLAM}}}\triangleq[p_{k}^{\top}\;q_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$, where $p_{k}\in\mathbb{R}^{3}$ denotes the position, $q_{k}\in S^{3}$ denotes the orientation in quaternion representation, and $\eta_{k}\in\mathbb{R}^{3+m}$ denotes the coefficients of the approximated GP magnetic field model.
@@
3.2 The Magnetic-Field-Aided Inertial Navigation System
MAINS proposed in [11] uses a low-cost IMU and a magnetometer array to provide low-drift inertial navigation. The state vector of MAINS at time step $k$ is defined as